OpenStudy (anonymous):

The length of the segment between the points (2a,a-4) and (4,-1) is 2 \sqrt{10} units. What is the product of all possible values for a ?

OpenStudy (anonymous):

(2a - 4)^2 + (a - 4 -(-1))^2 = (2a - 4)^2 + (a - 3)^2 = (4a^2 -16a + 16 + a^2 - 6a + 9 = 5a^2 - 22a + 25 sqrt(5a^2 - 22a + 25) = 2 sqrt(10) pm (5a^2 - 22a +25) = 40

OpenStudy (anonymous):

5a^2 - 22a - 15 = 0 and -5a^2 + 22a - 65 = 0

OpenStudy (anonymous):

40*0=0 the answer is 0?

OpenStudy (anonymous):

nope not!

OpenStudy (anonymous):

using distance formula, (2a-4)^2 + (a-4+1)^2 = sq(2sqrt(10)) 4a^2 -16a +16 +a^2 -6a +9 = 40 5a^2 -22a + 25 = 40 5a^2 - 22a -15 =0 5a^2 - 25a + 3a -15 =0 5a(a-5) + 3(a-5) = 0 (5a + 3) ( a-5) = 0 a = -3/5 or 5 so product = -3

OpenStudy (anonymous):

i was late coz i did some mistake in addition previously if you have noticed that

OpenStudy (anonymous):

are u still confused with this solution